Part a: You will find how to measure CPI-U U.S. city average by expenditure category. The two periods are September 2018 and September 2019. By what percent did this index change? Remember to show all work, starting with the formula and working through to the answer.

part b: Using 1982-84 as the base period, find the percentage change in the index from the base period to Q3 2019. Again, start with formula and derive answer typing everything.
Answer #1

Solution A

CPI U in US Expenditure Category for All Cities Average Unadjusted as per BLS website, is

September 2018= 252.439

September 2019= 256. 759

Thus percentage change = ((September's 2019- September 2018)/September 2018))×100

= ((256.759-252.439)/252.439)) × 100= 4.32×100/ 252.439= 1.7%.

Solution B

CPI U in US Expenditure Category for All Cities Average Unadjusted as per BLS website, is

Base Year 1982-84= 100.000

September 2019= 256. 759

Thus percentage change = ((September's 2019 - Base Year 1982 to 1984 /Base Year)×100

= ((256.759-100)/100)) × 100= 156.759%

In other words CPI U in US Expenditure Category for all cities has changed 1.56759 times from 1982-84 to 2019 September.